Linux Consultants HOWTO

Martin Michlmayr, tbm@sypher.com, http://www.sypher.com/tbm v1.69, 25 September 1996 This document contains a listing of companies providing commercial Linux related support. It is maintained by Martin Michlmayr <tbm@sypher.com>.

Contents

1 About this document 3

2 Copyright Information 3

3 America (North and South) 4 3.1 Rahim Azizarab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2 Scott Barker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 BRT Technical Services Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.4 Cabala Systems Corp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.5 M. Cooper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.6 Couvares Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.7 Crynwr Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.8 Datasoft Communications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.9 DigiTech, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.10 Dynamicro Consulting Limited . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.11 Electronic Information Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.12 Etc Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.13 Evolve Computer Solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.14 Christopher J. Fearnley . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.15 Global Proximity Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.16 G.M. Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.17 Granite Computing Solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.18 Steve Hedlund . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.19 Miguel de Icaza . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.20 Interactive Electronic Design Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.21 Internet Robotics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.22 Internet Systems and Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.23 Kensho, LLC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.24 Kracked Rock Komputing, LLC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.25 Kymsoft Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.26 NET-Community . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.27 The Network Wizards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.28 Practical Network Design . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.29 PSA AnswerDesk – Pencom Systems Administration . . . . . . . . . . . . . . . . . . . . . . . 23 3.30 RCT Design . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.31 Silver Technologies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.32 Softcraft Impresa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.33 Earl A. Stutes, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.34 Triangle Software Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.35 Vennerable Consultants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.36 Vertex Development . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.37 Gregg Weber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.38 Worldwide Solutions, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.39 xprt Computer Consuting, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

4 Asia 31 4.1 Vision Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

5 Australia 32 5.1 Interactive Infotainment Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

6 Europe 32 6.1 Bitbybit Information Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 6.2 Cistron Internet Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 6.3 Double Barrel Consultancy and Productions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 6.4 EDV-Beratung Friedrich Ganter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 6.5 Gärtner Datensysteme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 6.6 Ing. Büro Gerichhausen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 6.7 Guide Datakonsult . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 6.8 Information Cascade Ltd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 6.9 it-analyse as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 6.10 Laboratory for Micro Enterprise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 6.11 LEMIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.12 Netsol Network Solutions Oy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 6.13 B. Neubert EDV-Dienstleistungen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 6.14 NorConnect Internet Services AS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6.15 noris network GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 6.16 PROFUSO di Zanetti Giuseppe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 6.17 Signum Support AB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 6.18 Softmedia Spa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6.19 Tarkus AB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

1 About this document

This is the Linux Consultants HOWTO. It is a listing of companies providing commercial Linux related support. It is maintained by Martin Michlmayr <tbm@sypher.com>. If you need to know more about the Linux Documentation Project or about Linux HOWTO’s, feel free to contact the supervisor Greg Hankins <gregh@sunsite.unc.edu>. Greg Hankins will post the listing to several national and international newsgroups on a monthly basis. In addition, the Consultants HOWTO is available via FTP at ftp://ftp.sypher.com/tbm/HOWTO or can be found on the World Wide Web at http://www.sypher.com/tbm/Consultants-HOWTO. Companies providing Linux support are invited to fill out the following form and contact me at tbm@sypher.com. If you send me an entry for your product, please don’t forget to send me timely updates if appropriate.

     Name:
     Address:
     Phone:
     Fax:
     EMail:
     Url:
     Contact:
     Type of support (via phone, EMail, inhouse, training etc.):
     Sample prices:

2 Copyright Information

This HOWTO is Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996 by Martin Michlmayr.

A verbatim copy may be reproduced or distributed in any medium physical or electronic without permission of the author. Translations are similarly permitted without express permission if it includes a notice on who translated it.

Short quotes may be used without prior consent by the author. Derivative work and partial distributions of the Consultants HOWTO must be accompanied with either a verbatim copy of this file or a pointer to the verbatim copy.

Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

We further want that all information provided in the HOWTOs is disseminated. If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at gregh@sunsite.unc.edu.

3 America (North and South)

3.1 Rahim Azizarab

          1925 S 3rd Ave
          Maywood,  IL  60153
          U.S.A.

Phone:

708-344-6994

Fax:

None

EMail:

rahim@megsinet.net or ray@tezcat.com

Url:

None

Contact:

Rahim Azizarab

Type of support:

Phone, EMail, inhouse, training.

Sample prices:

Free for individuals.

Last modified:

September 18, 1996.

3.2 Scott Barker

     Calgary
     Alberta
     Canada

Phone:

(403) 209-9406 Fax:

(403) 285-1399 EMail:

scott@galileo.cuug.ab.ca Url:

http://www.cuug.ab.ca:8001/˜ barkers Contact:

Scott Barker Type of support:

I do everything, from vanilla installs, to specific customizations and hacks, to ongoing system administration. However, I do not do telephone-only support. I will do telephone support only if I can have direct access to the system, such as through the Internet or through a no-charge dial-in.

Sample prices:

$20/hour (personal); $30/hour (commercial). Last modified:

September 10, 1996.

3.3 BRT Technical Services Corporation

     50 Newtown Road
     Danbury, CT 06810
     U.S.A.

Phone:

(203) 748-5100 Fax:

(203) 792-0280 EMail:

info@brttech.com Url:

http://www.brttech.com Contact:

Chad Robinson Type of support:

Phone, EMail, inhouse, training. Sample prices:

Negotiated with client. Last modified:

September 20, 1996.

3.4 Cabala Systems Corp.

          801 3760 Albert St.,
          Burnaby, B.C.
          V5C 5Y8
          Canada

Phone:

(604) 294-1721 Fax:

(604) 294-1751 EMail:

kcarson@direct.ca Url:

None Contact:

Kevin Carson Type of support:

Phone, EMail, onsite, administration, training and programming. We cover all Linux related areas but specialities comprise of X11 graphics, security, networking, and 24/7/365 applications.

Sample prices:

Dependent on activity/contract. Last modified:

September 23, 1996.

3.5 M. Cooper

          PO Box 237
          St. David, AZ 85630
          U.S.A.

Phone:

(520) 720-9431 Fax:

None EMail:

thegrendel@theriver.com Url:

http://personal.riverusers.com/˜ thegrendel Contact:

M. Leo Cooper Type of support:

Phone, EMail. We offer set-up, programming and training. Sample prices:

US$20-$50/hour, negotiable. Last modified:

September 20, 1996.

3.6 Couvares Consulting

          211 West Gilman St.
          Suite 3W
          Madison, WI 53703
          U.S.A.

Phone:

+1 (608) 256-6901 Fax:

None EMail:

couvares@family.hampshire.edu Url:

None Contact:

Peter Couvares Type of support:

We offer phone/EMail support, installation, ongoing administration, training, programming, and specialized consulting services for Free Software, Linux, and other Unix environments.

Sample prices:

US$50/hour (commercial), US$40/hour (non-profit); sliding scale for personal. Last modified:

September 9, 1996.

3.7 Crynwr Software

          521 Pleasant Valley Rd.
          Potsdam, NY 13676
          U.S.A.

Phone:

+1-315-268-1925 Fax:

+1-315-268-9201 EMail:

info@crynwr.com Url:

http://www.crynwr.com Contact:

Russell Nelson Type of support:

Qualified to do any kind of support; specializing in networking drivers. Sample prices:

US$100/hour; fixed-price quotes for your task. Last modified:

September 21, 1996.

3.8 Datasoft Communications

          62 Rock Fernway
          Willowdale, ON M2J 4N5
          Canada

Phone:

416-756-4497 Fax:

416-756-3301 EMail:

mark@datasoft.com Url:

http://www.datasoft.com Contact:

Mark Buckaway Type of support:

Phone, EMail, inhouse; also setup and installation consulting, web page hosting. Visit our web page for details.

Sample prices:

CAN$85/hour (1 hour minimum) with flat rate contracts available. Last modified:

September 13, 1996.

3.9 DigiTech, Inc.

          301 N. Ankeny Blvd., Suite 100
          Ankeny, IA 50021
          U.S.A.

Phone:

515/964-8032 Fax:

515/965-7626 EMail:

digitech@ecity.net Url:

http://george.ecity.net/˜ digitech Contact:

Non specified Type of support:

Phone, EMail, inhouse, onsite, install, hardware, software, etc. Sample prices:

          Hardware/software support                     US\$52/hour
          Networking                                    US\$65/hour

Last modified:

September 15, 1996.

3.10 Dynamicro Consulting Limited

          282 Waratah Avenue
          Newmarket, Ont. L3Y 5J2
          Canada

Phone:

Available to clients. Fax:

Available to clients. EMail:

glouis@dynamicro.on.ca Url:

None Contact:

Greg Louis Type of support:

Phone, EMail, inhouse, training. Also custom software development for scientific, engineering, monitoring/control or business applications.

Sample prices:

CAN$125/hour for short-term contracts, negotiable for longer commitments. Last modified:

September 13, 1996.

3.11 Electronic Information Systems

     8758 Brook Road
     McLean, VA 22102
     U.S.A.

Phone:

(703)281-0409

Fax:

None

EMail:

ghurwitz@tjhsst.edu

Url:

http://bhurwitz.erols.com/SS

Contact:

Gus Hurwitz

Type of support:

EMail preferred, phone very doable, inhouse can be arranged, individual training with inhouse, group training available also.

Sample prices:

Simple questions via EMail are generally handled at no charge, ongoing EMail consulting US$25-$50, ongoing EMail support contract (seperate mail addresses available, are handeled first) available for about US$10 a month for 6 months or more. Phone support available if local to (703)281 exchange, or you pick up the tab, generally used as an extension to EMail support. Inhouse setup as little as US$50, with network setup about US$100, if hardware setup about US$150. The more systems we setup, the less it costs. Classes available for more than 5 people: topics ranging from simple installation, to C/C++ development, to system administration, to server setup and administration, etc. Students are generally welcome (and encouraged) to maintain EMail corespondance after the class, asking questions, and sharing experiences. If hourly rates are perfered, we like to stay below US$30, and are often willing to work for advertising space or endorsements.

Last modified:

September 20, 1996.

3.12 Etc Services

     31 N. Vivyen St.,
     Bergenfield, NJ 07621-1529
     U.S.A.

Phone:

(201) 385-7113 Fax:

None EMail:

swebster@carroll.com Url:

None Contact:

Scott D. Webster Type of support:

Inhouse. System setup and administration, TCP/IP network configuration, web server setup (not content creation), Linux LAN servers for DOS/Windows/Mac, DNS configuration, experience with Linux on x86, Linux on DEC Alpha (AXP), Digital Unix on Alpha (AXP), and Solaris on SPARC and UltraSPARC.

Sample prices:

Call to negotiate. Last modified:

September 15, 1996.

3.13 Evolve Computer Solutions

     1523 E. Valley Parkway, Suite 237
     Escondido, CA 92027
     U.S.A.

Phone:

(619)-746-2757 Fax:

(619)-751-0612 EMail:

evolve@sierra.net Url:

None Contact:

Sean P. Kane Type of support:

Phone, EMail, inhouse, training. Sample prices:

US$40/hour. Last modified:

September 17, 1996.

3.14 Christopher J. Fearnley

          4807 Chester AVE APT 207
          Philadelphia PA 19143-3441
          U.S.A.

Phone:

215-724-2265 Fax:

None EMail:

cjf@netaxs.com Url:

http://www.netaxs.com/˜ cjf Contact:

Chris Fearnley Type of support:

Phone, inhouse, Internet. Sample prices:

US$500/day. Last modified:

September 17, 1996.

3.15 Global Proximity Corporation

     959 Alice Street
     Woodstock
     Ontario N4S 8B1
     Canada

Phone:

519-421-3541 Fax:

519-421-2107 EMail:

CTyler@Oxford.net Url:

http://www.he.net/˜ globprox Contact:

Chris Tyler Type of support:

Full range of local and remote consulting services, ranging from installation support or the creation of small CGI scripts through the provision of full-service managed-Intranet services (including HW/SW/Comm/Distaster Recovery planning/Admin).

Sample prices:

Base price CAN$100/hour (about US$80), lower for large projects, higher for urgent work. Last modified:

September 12, 1996.

3.16 G.M. Systems

     2570 Seaview Road
     Mill Bay, BC. V0R 2P0
     Canada

Phone:

(604) 743-1500. Please note that the area code will change to 250 in October 1996. Fax:

(604) 743-1500 EMail:

george@gmsys.com Url:

None Contact:

George A. Farris Type of support:

Phone, EMail, inhouse, training, equipment sales and setup with Linux pre-installed. Sample prices:

Contract dependent. Last modified:

September 13, 1996.

3.17 Granite Computing Solutions

          P.O. Box 270103
          Fort Collins, CO 80527-0103
          U.S.A.

Phone:

+1 (970) 225-2370 Fax:

None EMail:

granite@SoftHome.Net Url:

http://www.SoftHome.Net/granite Contact:

Brian Grossman <brian@SoftHome.Net> Type of support:

In person, phone, EMail, Internet remote (telnet), training developed on request for Internet and Unix topics.

Sample prices:

US$45/hour for technical support. Last modified:

September 14, 1996.

3.18 Steve Hedlund

          Las Vegas, NV
          U.S.A.

Phone:

702-254-6749 Fax:

None EMail:

ssh@nevadacom.com Url:

http://nevadacom.com Contact:

Steve Hedlund Type of support:

Phone, EMail, inhouse, training. Sample prices:

US$25-$40/hour. Last modified:

September 12, 1996.

3.19 Miguel de Icaza

          Av Copilco 163-22-1003
          Mexico City
          Mexico

Phone:

Mexico-(915)-622-4680 Fax:

None EMail:

miguel.de.icaza@linux.org Url:

None Contact:

Miguel de Icaza Type of support:

Phone, EMail, training. Kernel hacking, driver developement, system administration, applications programming, bug fixing of publicaly available software, extension of freely redistributable software.

Sample prices:

US$30-$50/hour. Last modified:

September 15, 1996.

3.20 Interactive Electronic Design Inc.

          338 Westcourt Pl.
          Waterloo, ON, N2L 2R7
          Canada

Phone:

519-725-7353 Fax:

519-725-7353 EMail:

orders@ied.com Url:

None Contact:

Jan Vicherek Type of support:

Phone, EMail, inhouse, training, tutoring, Internet connections setup, custom programming. Sample prices:

CAN$20-$40/hour plus expenses. Last modified:

September 24, 1996.

3.21 Internet Robotics

          131 Willow Lane, Floor 2
          Oxford, OH 45056
          U.S.A.

Phone:

513.523.5028

Fax:

513.523.7501

EMail:

support@bill.iac.net

Url:

http://www.iac.net/˜ mikeh

Contact:

Mike Heins

Type of support:

Specializing in online shopping with MiniVend, the GPL shopping cart. Expertise in all aspects of Linux, WWW administration, shell and Perl scripting and network configuration. Performance tuning. Phone and EMail, will travel for training courses.

Sample prices:

          Expert technical support (uncontracted):      US\$100/hour
          Monthly retainer (includes first hour):       US\$50
          Contract technical support (with retainer):   US\$85/hour
          Contract shopping software design:            US\$85/hour
          Contract shopping site HTML design:           US\$50/hour
          Configuration support:                        US\$50/hour

Last modified:

September 12, 1996.

3.22 Internet Systems and Services

          13628 Bent Tree Circle, Suite 201
          Centreville, VA 22020
          U.S.A.

Phone:

(703) 222-4243 Fax:

(703) 222-7320 EMail:

bass@silkroad.com Url:

http://www.silkroad.com Contact:

Tim Bass, Principal Systems Engineer Type of support:

Phone, EMail, inhouse. Sample prices:

US$85.00/hour plus travel expenses (20 hours normal minimum). Last modified:

September 15, 1996.

3.23 Kensho, LLC

          9444 Balboa Avenue Suite 235
          San Diego, California 92123
          U.S.A.

Phone:

+1 (619) 514-3939 Fax:

+1 (619) 514-3938 EMail:

info@kensho.com Url:

http://www.kensho.com Contact:

Sean Dreilinger <sean@kensho.com> Type of support:

Research, training, and marketing of networked information services, with an eye towards PD and GPL solutions, i.e. Linux!

Sample prices:

Free consultation. Per project and hourly rates negotiable. Last modified:

September 12, 1996.

3.24 Kracked Rock Komputing, LLC

     13 Fleetwood Drive
     Danbury, CT 06810-7009
     U.S.A.

Phone:

+1 203 778 6269 Fax:

+1 203 778 6292 EMail:

ken@eci.com Url:

None Contact:

Kenneth E. Nawyn Type of support:

Phone, EMail, inhouse systems administration, system configuration and tech support for Linux, FreeBSD, SunOS, Solaris, IRIX and AIX. Over ten years of Unix systems adminstration experience along with Perl and C coding experience. For more detailed description of our experience and services, please contact us.

Sample prices:

Base rate US$100/hour. For a more detailed rate schedule, please contact us. Last modified:

September 24, 1996.

3.25 Kymsoft Consulting

     745 Foothill Dr
     Eugene, OR 97405
     U.S.A.

Phone:

541-465-9008 Fax:

541-465-9052 EMail:

rick@efn.org Url:

http://www.efn.org/˜ rick Contact:

Rick Bronson Type of support:

Phone, EMail, inhouse, training (12 years Unix experience). Sample prices:

US$65/hour. Last modified:

September 12, 1996.

3.26 NET-Community

          38403 Pelton Road
          Willoughby, OH 44094
          U.S.A.

Phone:

1-800-919-0060 Fax:

1-216-953-5829 EMail:

sales@net-community.com Url:

http://www.net-community.com Contact:

Scott Christley Type of support:

Phone, EMail. NET-Community specializes in support for GNUstep and MediaBook. Our web site has details regarding the support plans.

Sample prices:

          US\$50/year for GNUstep Developer - EMail only
          US\$200/year for GNUstep Corporate - EMail only
          US\$2500/year for MediaBook Comprehensive - EMail and phone

Last modified:

September 20, 1996.

3.27 The Network Wizards

          2275 West Birdsong Rd
          West Valley City, UT 84119
          U.S.A.

Phone:

(801) 973-4170 Fax:

(801) 975-7355 EMail:

tnw@ix.netcom.net Url:

None Contact:

Pat Eyler <pate@gnu.ai.mit.edu> Type of support:

Phone or EMail (most things); inhouse, training, contract telework available. Sample prices:

Interesting work and worthy charities are free (my definition of worthy). Others average US$50/hour (corporate) or US$25/hour (personal).

Last modified:

September 12, 1996.

3.28 Practical Network Design

     9 Chambers Rd
     Mansfield, OH 44906-1302
     U.S.A.

Phone:

(419) 529-3841 Fax:

(419) 529-3625 EMail:

stick@richnet.net Url:

None Contact:

Chuck Stickelman Type of support:

Inhouse, phone, EMail. Can provide system and network planning, installation, support and administration. Training is also available for hardware service/support, system/network administration, and various user-level applications.

Sample prices:

Phone support is US$50/hour (1 hour minimum). Inhouse support can be contracted on an hourly basis at US$100/hour or by day/half-day or by project.

Last modified:

September 23, 1996.

3.29 PSA AnswerDesk – Pencom Systems Administration

     9050 Capital of Texas Hwy North
     Suite 310
     Austin, TX 78759
     U.S.A.

Phone:

(512) 343-1111 or contact your nearest Pencom office. Fax:

(512) 346-6444 EMail:

answer-info@pencom.com Url:

http://www.pencom.com/psa/answer.html Contact:

Randy Kirchhof Type of support:

AnswerDesk is an Open Systems phone support desk specifically designed to provide a backstop with difficult, senior administration issues. We do not provide end user support. We support administrators working with all Unix platforms, including Linux. Our client list includes a wide variety of Fortune 500 companies utilizing open systems, many with mission critical support requirements. Our technical database is supported by 200 of the best systems administrators in the business.

Sample prices:

AnswerDesk charges US$100/hour, with a one hour minimum. If we don’t solve your problem, you are not charged. Please request a brochure for more in-depth info.

Last modified:

September 24, 1996.

3.30 RCT Design

Phone:

805-683-9717 Fax:

805-683-0717 EMail:

bob@zooid.com Url:

http://www.zooid.com Contact:

Bob Tellefson Type of support:

Turnkey Linux Systems with support, Internet servers, web site design and cgi programming, remote system maintenance and/or training by phone or EMail.

Sample prices:

US$45/hour. Last modified:

September 14, 1996.

3.31 Silver Technologies

          Box 22123
          Regina
          Saskatchewan S4S 7H4
          Canada

Phone:

306-596-1332 Fax:

306-585-2878 EMail:

rcp@dlcwest.com Url:

None Contact:

Richard Perrin Type of support:

System and network administration, on-site consultation, programming, Linux publications. Sample prices:

$20-$50/hour CDN. Last modified:

September 12, 1996.

3.32 Softcraft Impresa

          2923 Cohansey Drive
          San Jose, California 95132
          U.S.A.

Phone:

+1.408.251.9820 Fax:

None EMail:

mauro@softcraft.com Url:

http://www.softcraft.com Contact:

Mauro DePalma Type of support:

We can fine tune homogeneous and heterogeneous networks with Linux/Unix nodes and others TCP/IP capable. We also have an evolving secure operating environment based on Linux which is complete but does not (yet) come with installation scripts/programs and is currently only available for Intel. We’ll call it Linux generations. Gladly demo it and use it to build a LAN/WAN with shadow/NYS/DNS for secure nodes, gateways and networks.

Sample prices:

US$90/hour. Really depends on the kind of work and duration. Last modified:

September 20, 1996.

3.33 Earl A. Stutes, Inc.

     P.O. Box 35567
     Monte Sereno, CA 95030-0567
     U.S.A.

Phone:

408.448.1089 Fax:

408.264.5129 EMail:

estutes@eas.com Url:

http://www.eas.com Contact:

Earl Stutes Type of support:

Phone, EMail, inhouse, training. Sample prices:

Upon Request. Last modified:

September 12, 1996.

3.34 Triangle Software Corporation

          2375 Riverglenn Circle
          Dunwoody, GA 30338-5940
          U.S.A.

Phone:

770-455-9089 Fax:

Upon request. EMail:

triangle@america.net Url:

None Contact:

Ben Cooper Type of support:

Inhouse ANSI C++ training with labs using G++ (very well reviewed by students). Sample prices:

US$5000 plus travel expenses for a five-day C++ class with up to 15 students, taught at your site. Last modified:

September 12, 1996.

3.35 Vennerable Consultants

          7216 Lincoln Avenue, Suite 101
          El Cerrito, CA 94530 3240
          U.S.A.

Phone:

510 525 2989 Fax:

510 525 8072 EMail:

jason@vennerable.com Url:

http://www.vennerable.com Contact:

Jason Venner Type of support:

Phone, EMail, network and inhouse. Sample prices:

Variable by job and customer. Last modified:

September 10, 1996.

3.36 Vertex Development

          11 Crestview Court
          Potomac, MD 20854
          U.S.A.

Phone:

301-340-2162 Fax:

301-309-6968 EMail:

jeff@vertexdev.com Url:

http://www.vertexdev.com Contact:

Jeff Barr Type of support:

Porting Unix and non-Unix software to Linux. User Interface design, development, and review. Applications and Systems software development.

Sample prices:

Porting and development projects are typically handled on a firm-fixed price basis. Other services are supplied at US$120-$150/hour, depending on duration and travel requirements.

Last modified:

September 15, 1996.

3.37 Gregg Weber

     1076 Carol Lane \#3
     Lafayette, CA 94549
     U.S.A.

Phone:

510 283-6264 Fax:

None EMail:

gregg@netcom.com Url:

None Contact:

Gregg Weber Type of support:

On site in SF bay area, or can travel for big jobs. Sample prices:

US$55/hour. Last modified:

September 9, 1996.

3.38 Worldwide Solutions, Inc.

     4450 Arapahoe Ave., Suite 100
     Boulder, CO 80303
     U.S.A.

Phone:

+1.303.581.0800 Fax:

None EMail:

info@wwsi.com Url:

http://www.wwsi.com Contact:

Blaine Berger Type of support:

Phone, EMmail, Internet, on-site, training, installations, configuration, specification, architecture, engineering, integration, etc.

Sample prices:

Variable. Monthly support agreements or project based statement of work pricing preferred. Standard hourly rates available for those with specific requirements.

Last modified:

September 15, 1996.

3.39 xprt Computer Consuting, Inc.

     731 Voyager
     Houston, TX 77062
     U.S.A.

Phone:

713 480 8649 Fax:

713 480 7357 EMail:

jody@sccsi.com Url:

http://zeus.xprt.com/˜ info/xprt.html Contact:

Jody Winston Type of support:

Phone, EMail, training, custom software, custom device drivers, custom application development. Sample prices:

Either by the hour starting at US$100 or a firm fixed price. Last modified:

September 12, 1996.

4 Asia

4.1 Vision Systems

          179-C-1 Model Town
          Lahore-54700
          Pakistan

Phone:

(042) 583-6152 or (042) 586-9353 Fax:

92-42-5834920 EMail:

abnaseem@paknet1.ptc.pk Url:

http://www.freenet.edmonton.ab.ca/˜ abidn Contact:

Abid Nasim Type of support:

Training, configuration and BBS. Sample prices:

Upon request. Last modified:

September 18, 1996.

5 Australia

5.1 Interactive Infotainment Systems

          Sydney, NSW
          Australia

Phone:

015-955-042 Fax:

None EMail:

anthony@rumble.waratah.id.au Url:

http://www.infotainment.com.au Contact:

Anthony Rumble Type of support:

Phone, EMail, inhouse, training, porting, development, web/CGI, networking. Sample prices:

Upon request. Last modified:

September 24, 1996.

6 Europe

6.1 Bitbybit Information Systems

          Kluyverweg 2A
          2629 HT Delft
          The Netherlands

Phone:

+31.15.2682569 Fax:

+31.15.268.25.30 EMail:

info@bitbybit-is.nl Url:

http://www.bitbybit-is.nl Contact:

Erno Marks Type of support:

Installation, troubleshooting, training, Internet and Intranet connectivity. Sample prices:

US$60-$75/hour. Last modified:

September 13, 1996.

6.2 Cistron Internet Services

          PoBox 297
          2400 AG
          The Netherlands

Phone:

+31-172-419445 Fax:

+31-172-430979 EMail:

linux@cistron.nl Url:

http://www.cistron.nl Contact:

Danny ter Haar Type of support:

Author of large parts of Linux (sysvinit, minicom, network code and kernel code). Maintainer of many Debian GNU/Linux packages. Internet and Intranet provider including firewalls and leased lines.

Sample prices:

HFL 150/hour. Last modified:

September 24, 1996.

6.3 Double Barrel Consultancy and Productions

          Sportstraat 28

Phone:

32-9-2227764 Fax:

32-9-2227764 EMail:

mvergall@double-barrel.be Url:

http://www.double-barrel.be Contact:

Michael C. Vergallen Type of support:

Phone, EMail, inhouse. Sample prices:

1500 BEF/hour V.A.T. excl. (1 hour minimum). Last modified:

September 24, 1996.

6.4 EDV-Beratung Friedrich Ganter

          Grazerstra{\ss}e 26a
          A-8045 Graz
          Austria

Phone:

+43 316 688112 or +43 664 1021621 Fax:

None EMail:

ganter@fvkma.tu-graz.ac.at Url:

http://fvkma.tu-graz.ac.at/ganter/ganter.html Contact:

Friedrich Ganter Type of support:

Phone, EMail, inhouse, training. Sample prices:

US$30-$60/hour. Last modified:

September 12, 1996.

6.5 Gärtner Datensysteme

          Hamburger Str. 273a
          38114 Braunschweig
          Germany

Phone:

+49 531 2335555 Fax:

+49 531 2335556 EMail:

info@gaertner.de Url:

http://www.gaertner.de Contact:

Gärtner Datensysteme Type of support:

Phone, EMail, inhouse, training. Sample prices:

Depending on the job, starting at DM 1800 for a full day. Last modified:

September 13, 1996.

6.6 Ing. Büro Gerichhausen

          Steinsstra{\ss}e 103
          D-41199 M\"{o}nchengladbach
          Germany

Phone:

+49 2166 15068 Fax:

+49 2166 10159 EMail:

support@mail.niederrhein.de Url:

http://www.niederrhein.de Contact:

Christoph Adomeit Type of support:

Inhouse, training, phone, EMail, Internet. Sample prices:

DM 155/hour. Last modified:

September 12, 1996.

6.7 Guide Datakonsult

          Gustavslundsv. 12
          167 51 Bromma
          Sweden

Phone:

+46 70 745 52 39 Fax:

+46 634 46 00 EMail:

mge@guide.se Url:

http://www.guide.se Contact:

Mikael Geijer Type of support:

EMail, inhouse. Sample prices:

On discussion. Last modified:

September 15, 1996.

6.8 Information Cascade Ltd

     87 Burley Road
     Bransgore
     BH23 8BA
     UK

Phone:

0181-424-9691 Fax:

Upon request. EMail:

gps@trix.dircon.co.uk Url:

http://www.users.dircon.co.uk/˜ trix Contact:

Graham Swallow Type of support:

Linux lnstallation and configuration, SVR4 admin, site admin, network maintenence, Internet, servers and clients, shell scripts, document processing e.g. - Get your Quality Procedures Manual on-line (Your users too).

Sample prices:

£27/hour plus costs. Last modified:

September 22, 1996.

6.9 it-analyse as

          Breist{\o}lsveien 14
          N-5035 Bergen-Sandviken
          Norway

Phone:

+4755962221 Fax:

None EMail:

magnus@login.eunet.no Url:

None Contact:

Magnus Alvestad Type of support:

Phone, EMail, inhouse, training. Sample prices:

US$80-$100/hour. Last modified:

September 13, 1996.

6.10 Laboratory for Micro Enterprise

          Oakley Court
          Benson
          Wallingford, OX10 6QH
          UK

Phone:

(44) (0)1491 832932 Fax:

(44) (0)1491 832932 EMail:

lme@storrs.demon.co.uk Url:

http://www.i-way.co.uk/˜ storrs Contact:

John Storrs Type of support:

As required. Our focus is to support very small enterprises (most of us) with affordable systems and tools. See our web pages for details.

Sample prices:

By negotiation. Last modified:

September 15, 1996.

6.11 LEMIS

          Schellnhausen 2
          36325 Feldatal
          Germany

Phone:

+49-6637-919123 Fax:

+49-6637-919122 EMail:

lemis@lemis.de Url:

None Contact:

Greg Lehey Type of support:

Phone, EMail, fax, inhouse, training. Sample prices:

Typically US$100 or DM 150/hour. Significant reductions for non-profit organizations. Last modified:

September 23, 1996.

6.12 Netsol Network Solutions Oy

          Hiihtotie 3A
          FIN-01280 Vantaa
          Finland

Phone:

+358 50 5666898 Fax:

+358 9 34090404 EMail:

netsol@netsol.fi Url:

http://www.netsol.fi Contact:

Timo Virtaneva Type of support:

Inhouse, training. Support for Internet connections, firewalls and networking. Sample prices:

Upon request. Last modified:

September 15, 1996.

6.13 B. Neubert EDV-Dienstleistungen

          Parkweg 16
          Scheidt
          66133 Saarbr\"{u}cken
          Germany

Phone:

0681-817629 Fax:

Upon request. EMail:

neubert@lusi.uni-sb.de (bevorzugt/preferred)

Url:

http://www.uni-sb.de/˜ ph11rsbn Contact:

Boris Neubert Type of support (German):

Die Beratungstätigkeit umfaßt Hilfestellung bei Hardware- und Software-Kauf sowie Installation von Hardwarekomponenten und Anwenderprogrammen im DOS-, Windows- und Unix-Bereich, Netzwerklösungen im Unix-Bereich, Vermittlung von Netzzugängen, sowie die allgemeine Informationsbeschaffung im wissenschaftlichen Bereich und verwandte Dienstleistungen.

Type of support (English):

Consulting comprises assistance with hardware and software purchase, installation of hardware components and user programs under DOS, Windows and Unix, networking solutions under Unix, mediation of network access, general information retrieval in the scientific field and related services.

Sample prices:

DM 40/hour (zuzüglich Spesen/plus expenses). Last modified:

September 13, 1996.

6.14 NorConnect Internet Services AS

     Boks 9273 Gr\"{o}nland
     N-0134 Oslo
     Norway

Phone:

+47 22 17 41 70 Fax:

+47 22 17 64 99 EMail:

post@norconnect.no Url:

http://www.norconnect.no Contact:

Vidar Hokstad <vidarh@norconnect.no> Type of support:

Phone, EMail, inhouse, training, programming, planning, maintenance of systems, phone support for customers of our customer, porting of programs etc.

Sample prices:

Support agreements starting from 500,- NOK (approximately US$75), a month fixed fee (phone and EMail support, and reduced rates for other services). Projects or one time jobs starting at 650,- NOK (approximately US$110)/hour, with large discounts possible for larger projects, depending on the type of work, and whether our customer wants the project tied to one of our consultants, or not.

Last modified:

September 15, 1996.

6.15 noris network GmbH

     Schleiermacherstrasss 12
     90491 N\"{u}rnberg
     Germany

Phone:

+49 911 59818 0 Fax:

+49 911 59818 11 EMail:

info@noris.de Url:

http://www.noris.de Contact:

Matthias Urlichs <smurf@noris.de> Type of support:

Remote login, phone, EMail, inhouse. Sample prices:

DM 150/hour for phone or inhouse support. Last modified:

September 17, 1996.

6.16 PROFUSO di Zanetti Giuseppe

          via Vergani, 11
          I35031 Abano Terme (Padova)
          Italy

Phone:

++39 348 2220811 Fax:

None EMail:

profuso@campiello.it (will be profuso@profuso.com) Url:

http://www.campiello.it/PROFUSO (will be http://www.profuso.com) Contact:

Giuseppe Zanetti <beppe@iperv.it> Type of support:

All type of support for the region Veneto, EMail and phone for other regions. Sample prices:

Upon request. Last modified:

September 15, 1996.

6.17 Signum Support AB

          Teknikringen 8
          S - 583 30 Link\"{o}ping
          Sweden

Phone:

+46 (0)13 21 46 00 Fax:

+46 (0)13 21 47 00 EMail:

info@signum.se Url:

http://www.signum.se Contact:

Magnus Redin <redin@signum.se> Type of support:

Phone, fax, EMail and online. Specialised in the Readynet Internet server and the Readycom conferensing system, both are in house developed Free Software.

Sample prices:

Consulting US$110/hour, Readynet Internet server installation and support approximately US$4300. Last modified:

September 9, 1996.

6.18 Softmedia Spa

          via Pietrasanta 14
          Milano
          Italy

Phone:

(+39) 335-417677 Fax:

(+39) 2- 5730.7652 EMail:

venta@softmedia.it Url:

http://www.softmedia.it Contact:

Paolo Ventafridda Type of support:

Phone, inhouse, training and C programming. Sample prices:

Upon request. Last modified:

September 17, 1996.

6.19 Tarkus AB

          Skommartorpsv. 24
          735 38 Surahammar
          Sweden

Phone:

(+46) 70 7522246 Fax:

(+46) 220 33907 EMail:

tarkus@tarkus.se Url:

http://www.tarkus.se Contact:

Carlos Vidal, Bo Berggren Type of support:

Mostly via phone or EMail. Inhouse if required. We work in Scandinavia, Germany and Switzerland. Sample prices:

From 0 to 700 SEK/hour (US$0-$100) depending on the job. Many ”easy” questions are answered for free.

Last modified:

September 13, 1996.